OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (−6, 3). y = three halvesx − 1 y = three halvesx + 3 y = −two thirdsx − 1 y = −two thirdsx + 3

OpenStudy (anonymous):

help please!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

\[y = (3/2)x + 4\] the slope of the line is m = 3/2. since you are looking for the perpendicular line through (-6,3), you will need to take the negative reciprocal of the slope and use that as the slope of your perpendicular line. then use point slope formula \[(y-y _{1}) = m(x-x _{1})\] and plug in your values. you should get the following \[(y-3) = -(2/3)(x+6)\] solve for y to get your equation of the perpendicular line passing through (-6,3).